# A Wasserstein perspective of Vanilla GANs

### v1

@@ 4 Oracle inequality for Vanilla GAN in Wasserstein distance
Let $L>2$ and $B\in(0,\infty].$ For the empirical risk minimizer $\hat{G}_{n}$ from (10) with $\mathcal{W}=\operatorname*{Lip}(L,B)$ we have
@@
Note that the discriminator class $\operatorname*{Lip}(L,B)$ admits no finite dimensional parameterization and is therefore not feasible in practice. We will return to this issue in Section 5. The terms in (19) can be interpreted analogously to the interpretation of the bound in Lemma 2.1, but here we have an oracle inequality with respect to the Wasserstein distance. The first term is the approximation error. It is large when $\mathcal{G}$ is not flexible enough to provide a good approximation of $\mathbb{P}^{*}$ by $\mathbb{P}^{G}$ for some $G\in\mathcal{G}$. The second term refers to the stochastic error. With a growing number of observations the empirical measure $\mathbb{P}_{n}$ converges to $\mathbb{P}^{*}$ in Wasserstein distance, see Dudley, (1969), and thus the stochastic error converges to zero. Together with the bounds on $\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})$ by Schreuder, (2020) we conclude the following:
@@
Let $L>2,B\in(0,\infty]$. For some constant $c>0$ the empirical risk minimizer $\hat{G}_{n}$ from (10) with $\mathcal{W}=\operatorname*{Lip}(L,B)$ satisfies
$\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}(Z)})]\leq$ $\inf_{\begin{subarray}{c}G^{*}\colon\mathcal{Z}\rightarrow\mathcal{X}\end{subarray}}\Big{\{}c[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})]^{1;1/2}+c[\inf_{G\in\mathcal{G}}\|G-G^{*}\|^{1;1/2}_{\infty}]\Big{\}}$
$\qquad\qquad\qquad+c\begin{cases}n^{-1/2d},&d>2,\\ n^{-1/4}(\log n)^{1/2},&d=2,\\ n^{-1/4},&d=1.\end{cases}$
If there is some $G^{*}$ such that *$\mathbb{P}^{*}=\mathbb{P}^{G^{*}(Z)},$ which is commonly assumed in the GAN literature, see e.g. Stéphanovitch et al., (2023),  the first term vanishes and the approximation error is bounded by $\inf_{G\in\mathcal{G}}[\|G-G^{*}\|^{1;1/2}_{\infty}]$. In the bound of the stochastic error we observe the curse of dimensionality: For large dimensions $d$ the rate of convergence $n^{-1/2d}$ deteriorates.
To allow for a dimension reduction setting, we adopt the miss-specified setting from Vardanyan et al., (2023, Theorem 1). In this scenario we can conclude statistical guarantees for Vanilla GANs that are comparable to the results obtained for Wasserstein GANs by Schreuder et al., (2021, Theorem 2). Since Theorem 3.1 does not show equivalence of the Vanilla GAN distance and the Wasserstein distance, we lose a square root in the rate.
@@
Let $L>2,B\in(0,\infty]$ and $M>0$. The empirical risk minimizer $\hat{G}_{n}$ from (10) with $\mathcal{W}=\operatorname*{Lip}(L,B)$ satisfies
$\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}(Z)})]$ $\leq\inf_{\begin{subarray}{c}G^{*}\in\operatorname*{Lip}(M,\mathcal{Z})\end{subarray}}\Big{\{}c[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})]^{1;1/2}+c[\inf_{G\in\mathcal{G}}\|G-G^{*}\|^{1;1/2}_{\infty}]\Big{\}}$
$\qquad\qquad\qquad+c\begin{cases}n^{-1/2d^{*}},&d^{*}>2\\ n^{-1/4}(\log n)^{1/2},&d^{*}=2\\ n^{-1/4},&d^{*}=1,\end{cases}$
for some constant $c$ depending on the dimension $d^{*}$ of the latent space $\mathcal{Z}$ and $M$.
The Wasserstein distance $\mathsf{W}_{1}(\mathbb{P}^{G^{*}(Z)},\mathbb{P}^{*})$ now includes also the error due to the dimension reduction while the stochastic error is determined by the potentially much smaller dimension $d^{*}<d$ of the latent space. Compared to Corollary 4.2, the only price for this improvement is the additional Lipschitz restriction on $G^{*}$. We observe a trade-off in the choice of $d^{*}$, since large latent dimensions reduce the approximation error for $\mathbb{P}^{*}$, but increase the stochastic error term.
Theorem 4.3 reveals why Vanilla GANs do perform well in high dimensions. This phenomenon could not be explained in previous work on Vanilla GANs. Belomestny et al., (2021) and Biau et al., (2018) both obtain rates in the Jensen-Shannon distance.
@@ 5 Vanilla GANs with network discriminator
In practice, both $\mathcal{G}$ and $\mathcal{D}$ are sets of neural networks. Our conditions on the generator class $\mathcal{G}$ are compactness and good approximation properties of some $G^{*}$ which is chosen such that $\mathbb{P}^{G^{*}}$ mimics $\mathbb{P}^{*}$. Since neural networks have a finite number of weights, and the absolute value of those weights is typically bounded, the compactness assumption is usually satisfied and neural networks enjoy excellent approximation properties, cf. DeVore et al., (2021).
@@
Controlling the Lipschitz constant while preserving the approximation properties is an area of ongoing research and is far from trivial. Without further restrictions on the class of feedforward networks, the Lipschitz constant would be a term that depends exponentially on the size of the network, see Liang, (2017, Theorem 3.2). Bounding the Lipschitz constant of a neural network is a problem that arises naturally in the implementation of Wasserstein GANs. Arjovsky et al., (2017) use weight clipping to ensure Lipschitz continuity. Later, other approaches such as gradient penalization (see Gulrajani et al., (2017), which was further developed by Wei et al., (2018), Zhou et al., (2019)), Lipschitz penalization (Petzka et al.,, 2017), or spectral penalization (Miyato et al.,, 2018) were introduced and have achieved improved performance in practice.
@@
Let $L>2,B\in(0,\infty]$. The empirical risk minimizer $\hat{G}_{n}$ from (10) with $\mathcal{W}\subseteq\operatorname*{Lip}(L,B)$ satisfies for some constant $c>0$ that
@@
$\qquad\qquad\qquad+c\big{[}\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})\big{]}^{1;1/2}.$
To apply this result, we must ensure that the Lipschitz constant of a set of neural networks $\mathcal{W}$ is uniformly bounded by some constant $L$. Adding penalties to the objective function of the optimization problem does not guarantee a fixed bound on the Lipschitz constant. Approaches such as bounds on the spectral or row-sum norm of matrices in feedforward neural networks ensure a bound on the Lipschitz constant, but lead to a loss of expressiveness when considering ReLU networks, even in very simple cases such as the absolute value, see Huster et al., (2019) and Anil et al., (2019). On the other hand, Eckstein, (2020) has shown that one-layer $L$ Lipschitz networks are dense (with respect to the uniform norm) in the set of all $L$ Lipschitz functions on bounded domains. While this implies that the discriminant approximation error converges to zero for growing network architectures, the density statement does not lead to a rate of convergence that depends on the size of the network.
@@
Let $L>2,B>\log(3)$ and $\Gamma>\max(L,2B)$. The empirical risk minimizer $\hat{G}_{n}$ from (10) with $\mathcal{W}\subseteq\mathcal{H}^{\alpha}(\Gamma)$ satisfies for any $G^{*}\in\operatorname*{Lip}(M,\mathcal{Z})$
$\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}(Z)})]\leq$ $c\big{[}\inf_{G\in\mathcal{G}}\|G^{*}-G\|_{\infty}^{\alpha}\big{]}^{1;1/2}+c\big{[}\mathsf{W}_{1}(\mathbb{P^{*}},\mathbb{P}^{G^{*}(Z)})^{\alpha}\big{]}^{1;1/2}$
@@
$+c\begin{cases}n^{-\alpha/2d^{*}},&2\alpha<d^{*}\\ n^{-1/4}(\log n)^{1/2},&2\alpha=d^{*}\\ n^{-1/4},&2\alpha>d^{*}\end{cases}$
for some constant $c$ depending on $d^{*},M$ and $\Gamma$.
@@
[…]
Since there are many different reasons why controlling the Hölder constant of neural networks is interesting (with stability probably being the most prominent one), Theorem 5.3 is of interest on its own. Combining Theorem 5.2 and Theorem 5.3 with a standard approximation result for the generator approximation error, such as Yarotsky, (2017, Theorem 1), leads to a rate of convergence.
@@
For $0<\alpha<1,$ $\Gamma>5,$ $M>0,$ $d^{*}>2\alpha$ and $n>2^{\frac{2d^{*}}{\alpha}}$ choose $\mathcal{G}$ as the set of ReLU networks with at most $c\cdot(\log(n)+1)$ layers, $c\cdot n^{-\alpha/2}(\log(n)+1)$ neurons and $c\cdot n^{-\alpha/2}(\log(n)+1)$ nonzero weights and $\mathcal{W}^{\prime}$ as the set of ReLU networks with at most $1+c\cdot\log(n)$ layers and $c\cdot n^{\frac{\alpha}{2(1-\alpha)}}\log^{2}(n)$ nonzero weights and neurons, where $c$ is a constant depending on $d,d^{*},\Gamma$ and $\alpha$. Then the empirical risk minimizer $\hat{G}_{n}$ from (10) with $\mathcal{W}=\mathcal{W}^{\prime}\cap\mathcal{H}^{\alpha}(\Gamma)$ satisfies
@@
This corollary shows that Vanilla GANs with a Hölder regular discriminator class are theoretically advantageous. Since $\alpha$ can be chosen arbitrarily close to one, this reveals why a Lipschitz regularization as implemented for Wasserstein GANs also improves the Vanilla GAN. An empirical confirmation can be found in Zhou et al., (2019).

### v2

@@ 4 Oracle inequality for Vanilla GANs in Wasserstein distance
Let $L>2$ and $B\in[1,\infty].$ For the empirical risk minimizer $\hat{G}_{n}$ from (8) with $\mathcal{W}=\operatorname*{Lip}(L,B)$ we have
@@
Note that the discriminator class $\operatorname*{Lip}(L,B)$ admits no finite dimensional parameterization and is therefore not feasible in practice. We will return to this issue in Section 5. The terms in (11) can be interpreted analogously to the interpretation of the bound in Lemma 2.1, but here we have an oracle inequality with respect to the Wasserstein distance. The first term is the approximation error. It is large when $\mathcal{G}$ is not flexible enough to provide a good approximation of $\mathbb{P}^{*}$ by $\mathbb{P}^{G(Z)}$ for some $G\in\mathcal{G}$. The second term refers to the stochastic error. With a growing number of observations the empirical measure $\mathbb{P}_{n}$ converges to $\mathbb{P}^{*}$ in Wasserstein distance, see Dudley, (1969), and thus the stochastic error converges to zero. Together with the bounds on $\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})$ by Schreuder, (2020) we conclude the following:
@@
Let $L>2,B\in[1,\infty]$. The empirical risk minimizer $\hat{G}_{n}$ from (8) with $\mathcal{W}=\operatorname*{Lip}(L,B)$ satisfies for some constant $c>0$ depending on $d,p$ and $L$ that
$\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}(Z)})]\leq$ $\inf_{G^{*}\colon\mathcal{Z}\rightarrow\mathcal{X}}\Big{\{}c[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})]^{1;1/2}+c[\inf_{G\in\mathcal{G}}\|G-G^{*}\|_{\infty}]^{1;1/2}\Big{\}}$
$\qquad\qquad\qquad+c\begin{cases}n^{-1/2d},&d>2,\\ n^{-1/4}(\log n)^{1/2},&d=2,\\ n^{-1/4},&d=1,\end{cases}$
where the infimum is taken over all Borel measurable functions $G^{*}\colon\mathcal{Z}\rightarrow\mathcal{X}$.
If there is some $G^{*}$ such that $\mathbb{P}^{*}=\mathbb{P}^{G^{*}(Z)},$ which is commonly assumed in the GAN literature, see e.g. Stéphanovitch et al., (2023),  the first term vanishes and the approximation error is bounded by $\inf_{G\in\mathcal{G}}[\|G-G^{*}\|^{1;1/2}_{\infty}]$. In the bound of the stochastic error we observe the curse of dimensionality: For large dimensions $d$ the rate of convergence $n^{-1/2d}$ deteriorates.
To allow for a dimension reduction setting, we adopt the miss-specified setting from Vardanyan et al., (2023, Theorem 1). In this scenario we can conclude statistical guarantees for Vanilla GANs that are comparable to the results obtained for Wasserstein GANs by Schreuder et al., (2021, Theorem 2). In view of Theorem 3.1 we expect a slower rate of convergence compared to Wasserstein GANs.
@@
Let $L>2,B\in[1,\infty]$ and $M>0$. The empirical risk minimizer $\hat{G}_{n}$ from (8) with $\mathcal{W}=\operatorname*{Lip}(L,B)$ satisfies
$\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}(Z)})]$ $\leq\inf_{\begin{subarray}{c}G^{*}\in\operatorname*{Lip}(M,\mathcal{Z})\end{subarray}}\Big{\{}c[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})]^{1;1/2}+c[\inf_{G\in\mathcal{G}}\|G-G^{*}\|_{\infty}]^{1;1/2}\Big{\}}$
$\qquad\qquad\qquad+c\begin{cases}n^{-1/2d^{*}},&d^{*}>2,\\ n^{-1/4}(\log n)^{1/2},&d^{*}=2,\\ n^{-1/4},&d^{*}=1,\end{cases}$
for some constant $c$ depending $d^{*},d,p,L$ and $M$.
The Wasserstein distance $\mathsf{W}_{1}(\mathbb{P}^{G^{*}(Z)},\mathbb{P}^{*})$ now includes an error due to the dimension reduction while the stochastic error is determined by the potentially much smaller dimension $d^{*}<d$ of the latent space. Compared to Corollary 4.2, the only price for this improvement is the additional Lipschitz restriction on $G^{*}$. We observe a trade-off in the choice of $d^{*}$, since large latent dimensions reduce the approximation error for $\mathbb{P}^{*}$, but increase the stochastic error term. Additionally, there is a trade-off in $M$. A larger constant $M$ results in a smaller value of $W_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})$, but increases the constant $c$. If the unknown distribution $\mathbb{P}^{*}$ is supported on a lower dimensional subspace and there exists a $G^{*}\in\operatorname{Lip}(M,\mathcal{Z})$ such that $\mathbb{P}^{G^{*}(Z)}=\mathbb{P}^{*}$, then the rate of convergence is solely determined by the dimension $d^{*}$ of $\mathcal{Z}$. This is true for the smallest possible $d^{*}$ for which a perfect $G^{*}$ exists, as well as any $d^{**}$ larger than $d^{*}$.
In many applications, the smallest possible $d^{*}$ is unknown. Theorem 4.3 covers both over- and underestimation of the true dimension of the lower dimensional subspace. If the choice of $d^{*}$ is too small, then $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})$ might not converge to $0,$ but the stochastic error still converges with the smaller rate $d^{*}.$ If $d^{*}$ is selected to be larger than the dimension of the lower dimensional subspace, then there could be a $G^{*}\in\operatorname{Lip}(M,\mathcal{Z})$ such that $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})=0$, but the stochastic rates converges only with rate $d^{*}.$ In the special case that a function $G^{*}\in\operatorname{Lip}(M,\mathcal{Z})$ exists such that $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{G^{*}(Z)})=0$ the rate $n^{-1/2d^{*}}$ is slower to the rate $n^{-1/d^{*}}$ obtained for the Wasserstein GAN by Schreuder et al., (2021). This is in line with Theorem 3.2 and Example 3.3.
However, Theorem 4.3 reveals why Vanilla GANs do perform well in high dimensions in the setting of an unknown distribution on a lower dimensional manifold. This phenomenon could not be explained in previous work on Vanilla GANs. Puchkin et al., (2024) and Biau et al., (2020) both obtain rates in the Jensen-Shannon distance.
@@ 5 Vanilla GANs with network discriminator
In practice, both $\mathcal{G}$ and $\mathcal{D}$ are sets of neural networks. Our conditions on the generator class $\mathcal{G}$ are compactness and good approximation properties of some $G^{*}$ which is chosen such that $\mathbb{P}^{G^{*}(Z)}$ mimics $\mathbb{P}^{*}$. Since neural networks have a finite number of weights, and the absolute value of those weights is typically bounded, the compactness assumption is usually satisfied and neural networks enjoy excellent approximation properties, cf. DeVore et al., (2021).
@@
Controlling the Lipschitz constant while preserving the approximation properties is an area of ongoing research and is far from trivial. Without further restrictions on the class of feedforward networks, the Lipschitz constant would be a term that depends exponentially on the size of the network, see Liang, (2017, Theorem 3.2). Bounding the Lipschitz constant of a neural network is a problem that arises naturally in the implementation of Wasserstein GANs. Arjovsky et al., (2017) use weight clipping to ensure Lipschitz continuity. Later, other approaches such as gradient penalization (see Gulrajani et al., (2017), which was further developed by Wei et al., (2018), Zhou et al., (2019)), Lipschitz penalization (Petzka et al.,, 2018), or spectral penalization (Miyato et al.,, 2018) were introduced and have achieved improved performance in practice.
@@
Let $L>2,B\in[1,\infty]$. The empirical risk minimizer $\hat{G}_{n}$ from (8) with $\mathcal{W}\subseteq\operatorname*{Lip}(L,B)$ satisfies
@@
$\qquad\qquad\qquad+c\big{[}\mathsf{W}_{1}(\mathbb{P}_{n},\mathbb{P}^{*})\big{]}^{1;1/2},$
for some constant $c>0$ depending on $d,p$ and $L.$
The approximation error of the discriminator depends on the supremum norm bound $B$ of the functions in $\mathcal{W}$. While the statement remains true for $B=\infty$, when approximating the set $\operatorname{Lip}(L,B)$, this bound will be essential. To apply this result, we must ensure that the Lipschitz constant of a set of neural networks $\mathcal{W}$ is uniformly bounded by some constant $L$. Adding penalties to the objective function of the optimization problem does not guarantee a fixed bound on the Lipschitz constant. Approaches such as bounds on the spectral or row-sum norm of matrices in feedforward neural networks ensure a bound on the Lipschitz constant, but lead to a loss of expressiveness when considering ReLU networks, even in very simple cases such as the absolute value, see Huster et al., (2019) and Anil et al., (2019). On the other hand, Eckstein, (2020) has shown that one-layer $L$ Lipschitz networks are dense (with respect to the uniform norm) in the set of all $L$ Lipschitz functions on bounded domains. While this implies that the discriminant approximation error converges to zero for growing network architectures, the density statement does not lead to a rate of convergence that depends on the size of the network.
@@
Let $L>2,B\in[1,\infty),$ $\Gamma>\max(L,2B)$and $M>0.$ The empirical risk minimizer $\hat{G}_{n}$ from (8) with $\mathcal{W}\subseteq\mathcal{H}^{\alpha}(\Gamma)$ satisfies
$\mathbb{E}[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{G}_{n}(Z)})]\leq$ $\inf_{G^{*}\in\operatorname*{Lip}(M,\mathcal{Z})}\Big{\{}c\big{[}\inf_{G\in\mathcal{G}}\|G^{*}-G\|_{\infty}^{\alpha}\big{]}^{1;1/2}+c\big{[}\mathsf{W}_{1}(\mathbb{P^{*}},\mathbb{P}^{G^{*}(Z)})^{\alpha}\big{]}^{1;1/2}\Big{\}}$
@@
$+c\begin{cases}n^{-\alpha/2d^{*}},&2\alpha<d^{*},\\ n^{-1/4}(\log n)^{1/2},&2\alpha=d^{*},\\ n^{-1/4},&2\alpha>d^{*},\end{cases}$
for some constant $c$ depending on $d^{*},d,p,L,M$ and $\Gamma$.
The lower bound on the Hölder constant of the discriminator class $\mathcal{W}$ is not overly restrictive when employing neural networks for this function class. Since $c$ is increasing in $\Gamma$, it is advantageous to control the value of $\Gamma$.
@@
$\big{\|}\Phi_{\varepsilon}-f\big{\|}_{\infty}\leq\varepsilon\qquad\text{and}\qquad\Phi_{\varepsilon}\in\mathcal{H}^{\alpha}\big{(}\max(L,2B)+\varepsilon\big{)}.$
Since there are many different reasons why controlling the Hölder constant of neural networks is interesting (with stability probably being the most prominent one), Theorem 5.3 is of interest on its own. Combining Theorem 5.2 and Theorem 5.3 with a standard approximation result for the generator approximation error, such as Yarotsky, (2017, Theorem 1), leads to a rate of convergence. The networks in $\mathcal{G}$ approximating the function $G^{*}\in\operatorname{Lip}(M,\mathcal{Z})$ are only required to be measurable without any additional smoothness assumption.
@@
For $0<\alpha<1,$ $\Gamma>5,$ $M>0,$ $d^{*}>2\alpha$ and $n>2^{\frac{2d^{*}}{\alpha}}$ choose $\mathcal{G}$ as the set of ReLU networks with at most $\lceil c\cdot\log(n)\rceil$ layers, $\lceil c\cdot n\log(n)\rceil$ nonzero weights and $\lceil c\cdot n\log(n)\rceil$ neurons and $\mathcal{W}^{\prime}$ as the set of ReLU networks with at most $\lceil c\cdot\log(n)\rceil$ layers, $\lceil c\cdot n^{\frac{\alpha}{2(1-\alpha)}}\log^{2}(n)\rceil$ nonzero weights and $\lceil c\cdot n^{\frac{\alpha}{2(1-\alpha)}}\log^{2}(n)\rceil$ neurons, where $c$ is a constant depending on $d,d^{*},\Gamma,M$ and $\alpha$. Then the empirical risk minimizer $\hat{G}_{n}$ from (8) with $\mathcal{W}=\mathcal{W}^{\prime}\cap\mathcal{H}^{\alpha}(\Gamma)$ satisfies
@@
From Theorem 5.3 we know that the set $\mathcal{W}^{\prime}\cap\mathcal{H}^{\alpha}(\Gamma)$ of ReLU networks of finite width and depth is nonempty. In practice, this corresponds to a discriminator network with a controlled Hölder constant. On a bounded domain, any Lipschitz function is a Hölder function. Corollary 5.4 shows that Vanilla GANs with a Hölder regular discriminator class are theoretically advantageous. The Hölder parameter $\alpha$ can be chosen arbitrarily close to one. On the one hand this reveals why a Lipschitz regularization as implemented for Wasserstein GANs also improves the Vanilla GAN. An empirical confirmation can be found in Zhou et al., (2019) and Section 7. On the other hand the corollary then requires more neurons in the discriminator than in the generator class which coincides with common practice.
The width of the generator networks in Corollary 5.4 can be improved by replacing $G^{*}\in\operatorname{Lip}(M,\mathcal{Z})$ by with $G^{*}\in\operatorname{C}^{n-1}(\mathcal{Z}),n\in\mathbb{N},$ whose $(n-1)$-th derivative is Lipschitz continuous with Lipschitz constant $M.$ Once more, this results in a trade-off, as $\big{[}\mathsf{W}_{1}(\mathbb{P^{*}},\mathbb{P}^{G^{*}(Z)})^{\alpha}\big{]}^{1;1/2}$ increases when $G^{*}$ is selected from a smaller set of functions.